OpenStudy (anonymous):

Alexander's hobby is dirt biking. On one occasion last weekend, he accelerated from rest to 17.8 m/s in 1.56 seconds. He then maintained this speed for 9.47 seconds. Seeing a coyote cross the trail ahead of him, he abruptly stops in 2.79 seconds. Determine Alexander's average speed for this motion. @Vaidehi09

OpenStudy (anonymous):

So the issue i'm having, is the last part, where u said vf=vo-at...am i finding vf in this or acceleration?

OpenStudy (anonymous):

i will repost the answer u posted before.

OpenStudy (anonymous):

we know that avg speed = (total distance traveled) / (total time take).........(i) now total time = ( 1.56 + 9.47 + 2.79 ) secs. the total distance covered is calculated as follows: the whole journey can be divided into three parts: 1) when he accelerated from rest 2) when he moves with constant speed 3) when he decelerated and came to a stop. now, find the distance covered in each of these parts using the kinematic equations. 1) given: u1 =0, a1 = 17.8m/s^2, t1 = 1.56 sec s1 = u1t1 + 1/2 a1 t1^2 s1 = 0 + 1/2 (17.8) (1.56)^2 s1 = 8.9 (1.56)^2 m also, the constant speed acquired v1 = u1 + a1 t1 v1 = (17.8) (1.56) m/s 2) given: constant speed v1 = (17.8)(1.56) m/s, t2 = 9.47 sec s2 = v1 * t2 s2 = (17.8) (1.56) (9.47) m 3) given: v3 = 0, t3 = 2.79 s, v1 = u3 = (17.8)(1.56) m/s first find the deceleration [-a3] v3 = u3 - a3 t3 u3 = a3 t3 a3 = u3/ t3 a3 = (17.8) (1.56) / (2.79) m/s^2 (v3)^2 = (u3)^2 - 2 a3 s3 (u3)^2 = 2 a3 s3 s3 = (u3)^2 / 2 a3 s3 = (17.8)^2 (1.56)^2 (2.79) / 2 (17.8) (1.56) s3 = (17.8) (1.56) (2.79) / 2 s3 = 8.9 (1.56) (2.79) m so, total distance covered = s1 + s2 + s3. substitute this in (i) and u'll get the avg speed. The Changes: 1) given: u1 = 0, v1 = 17.8, t1 = 1.56 use v1 = u1 + a1 t1.......to find a1. then sub that s1 = 1/2 a1 t1^2......to get s1. 2) given: constant speed = v1 = 17.8, t2 = 9.47 calculate s2 using, s2 = v1 t2 3) here, proceed the original way. the only change will be, v1 = u3 = 17.8

OpenStudy (anonymous):

u already know that Vf = 0. and Vo is given. so first find out a.

OpenStudy (anonymous):

6.38m/s^2...so this would be substituted into d=vot+1/2at^2?

OpenStudy (anonymous):

yeah, u can do that. but if u use v^2 = u^2 + 2as, it'll be easier since v = 0.

OpenStudy (anonymous):

ok so i did the calculations and i added up all the distances. 1193.18meters, which i have to divide by 13.82, which is the total time. are the numbers correct?

OpenStudy (anonymous):

oh, i didn't calculate all that. give me a sec.

OpenStudy (anonymous):

i got the total distance = 207.28m

OpenStudy (anonymous):

I got that before, but the answer on the website is 23.4m/s...if you divide 207.28 by 13.82, you get around 15m/s. http://www.physicsclassroom.com/calcpad/1dkin/problems.cfm number 32.

OpenStudy (anonymous):

u heard the guided solution? its not opening on my comp.

OpenStudy (anonymous):

the plugin doesn't exist anymore...so yea.

OpenStudy (anonymous):

btw, how did u get 1193.18 ?

OpenStudy (anonymous):

umm. i got 1010.72m for the last one with: d=Vo^2/2a d=17.8^2/2(6.38) d=1010.72m. Then i added the rest.

OpenStudy (anonymous):

oh never mind, my calculator did it wrong.

OpenStudy (anonymous):

24.83m

OpenStudy (anonymous):

Yea so using that i get 207.29m

OpenStudy (anonymous):

ah! got it. to get 24 as answer....proceed the the way i showed in the beginning. don't see the correction part. coz in the question on that site, 17.8 is given as acceleration and not velocity.

OpenStudy (anonymous):

ohhhhhhh i seeee. ok so i was correct the first time around...lol i'm so dumb. THANKS A TON! :)

OpenStudy (anonymous):

lol! sure, not a prob! :)